OpenStudy (comm.dan):

This is a clarification question Am I doing this right so far? This is Graphing Quadratic Functions First, you find the Axis Symmetry by using x=-b over 2a Second, you put the x back into the original equation Third, put the x and y together to make the vertex Fourth, the C is the y-intercept, and then rewrite the y-intercept as 0 (zero), then the y-intercept Then, graph the axis of symmetry, vertex, y-intercept, and the two other points

OpenStudy (comm.dan):

Here is an example of a problem \[y =x ^{2}-2\]

OpenStudy (comm.dan):

then after the x squared minus 2 there is minus 3

OpenStudy (comm.dan):

@seiga

OpenStudy (comm.dan):

Ok then. I will teach you how I was taught to do it. @seiga

OpenStudy (comm.dan):

I am upload the pages right now @Pologirl19

OpenStudy (comm.dan):

Thanks again for trying to help me @e.cociuba

OpenStudy (comm.dan):

y=-4x squared +12x-5

OpenStudy (comm.dan):

@Pologirl19

OpenStudy (comm.dan):

\[x =\frac{ -12}{ 2 (-4) } = \frac{ -12 }{ -8}=\frac{ 3 }{ 2 }=1.5 \]

OpenStudy (comm.dan):

\[y =-4(1.5^{2})+12(1.5)-5=-4(2.25)+18-5=-10+18-5=8-5=3\]

OpenStudy (comm.dan):

2 minutes ago I did the first step A few moments ago I did the second step

OpenStudy (comm.dan):

\[(1.5,3) Vertex y-intercept Rewrite y-intercept as (0,-5)\]

OpenStudy (comm.dan):

\[Axis of Symmetry is 1.5\]

OpenStudy (comm.dan):

I will do the graph now

OpenStudy (anonymous):

The 5th one where you plugging it in you said it was -10+18-5. It is -9+18-5. So then it would equal 4. not 3

OpenStudy (comm.dan):

-4*2.25=10.00 is what I got

OpenStudy (comm.dan):

My bad @Pologirl19

OpenStudy (comm.dan):

Thank you

OpenStudy (anonymous):

K well everything I see is good so you are good

OpenStudy (comm.dan):

Thank you. @Pologirl19 Do you have any questions?

OpenStudy (anonymous):

No I don't it was a clear to me

OpenStudy (comm.dan):

Ok @Pologirl19

OpenStudy (e.cociuba):

np:)